(a) worked out incorrectly; or (b) worked out correctly but using a method not specified in the problem statement; or (c) not3) Find the sum of Un=3 510 () (5 points)
Answer #1

54 2 t -Р n=3 4 4 tot f2 4 2 + 23 + 9 F 2 1+2+2+2 +2 27 az1 N=8 a (2-1) Sn = r So = 1. (2001) P-1 So = 28 1 - 256-1=255